Session: May 2001 (session 2)
Wavelength: 6.0 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Tonino Pisanu at Wed Jun 6 05:57:15 METDST 2001) Success
Success without any problem.

Effelsberg (A. Kraus at Wed Jun 6 09:02:40 METDST 2001) Minor failures
Due to problems with the telescope control computer, we lost data between UT2051-2104 and UT2219-2222.

Torun (K.Borkowski at Wed Jun 6 09:15:25 METDST 2001) Minor failures
BBC5 and BBC8 out of order (low levels). The scan at 00:06 and half of the next one lost were due to slewing through 360 deg in azimuth.

Noto (C.Trigilio at Wed Jun 6 10:51:45 METDST 2001) Success

Shanghai (Huang Xinyong at Thu Jun 7 02:38:30 METDST 2001) Success

Jodrell_Bank (Alastair Gunn at Thu Jun 7 11:25:20 METDST 2001) Success
No known problems.

Cambridge (Alastair Gunn at Thu Jun 7 11:25:44 METDST 2001) Success
No known problems.

Urumqi (Liu Xiang at Thu Jun 7 11:47:13 METDST 2001) Failure
No thin tape available.

Onsala (John Conway at Thu Jun 7 11:47:34 METDST 2001) Minor failures
Some passes shows high parity errors, possibility related to vacumn changes and tape shifts, continuing to invstigate.

Westerbork (Tony Foley at Mon Jun 11 10:46:51 METDST 2001) Minor failures
Lots of array sync errors from 18:00 to 20:00 may give some decorrelation